Twice a number,decreased by 9, is half the same number increased by 6
Answer by checkley77(12844)

2X-9=X/2+6
2X-.5X=6+9
1.5X=15
X=15/1.5
X=10 ANS.
PROOF:
2*10-9=10/2+6
20-9=5+6
11=11